OpenStudy (anonymous):

Someone Helpppp???!!! Step by step to solve? How many solutions does it have??? x-4y=12 and 5x-20y=60

OpenStudy (anonymous):

First x - 4y = 12: Get Y on one side, X on the other. to do that, we add 4y to both sides. x - 4y + 4y = 12 + 4y now we have x = 12 + 4y and we can't simplify it anymore, so we go to the other one. 5x-20y=60 Same as before, but now we want to put the Y value alone, so we subtract both sides by 5x 5x -20y - 5x = 60 -5x and we get. -20y = 60 -5x Now lets get rid of the 20 on the y, by dividing by 20 and we get: -20y/20 = (60 -5x)/20 which is: -y = 3 - 1/4x Not fully simplified because of the minus behind the y, so lets multiply it all by -1 and get rid of it: y = -3 + 1/4x

OpenStudy (anonymous):

Do you know what to do next?

OpenStudy (anonymous):

no honestly not .. im really bad at math . i have to look at what u typed so i can write it down

OpenStudy (anonymous):

so wait what is next

OpenStudy (anonymous):

i just read the msg

OpenStudy (anonymous):

see your private messages, i'll explain

OpenStudy (triciaal):

let's use substitution from the 1st eqn x = 12 + 4y put this in the 2nd eqn 5x - 20y = 60 5 (12 + 4y) - 20y = 60 to graph eqn 1 when x = 0 what is y when y = 0 what is x (0, -3) and (12, 0) to graph eqn 2 (0, -5) and (12, 0) find where they meet only 1 solution

OpenStudy (anonymous):

do u have algebra 2?